0 0
Add a comment Improve this question Transcribed image text
Answer #1

P = 120 W ; R = 20 ohms

part A

P = V2/R

or, emf = V =sqrt (RP) = 48.99 Volts

part B

V = IR

or, I = emf /R = 48.99/20 = 2.45 A